# System-Theoretic Analysis of Dynamic Generalized Nash Equilibrium Problems – Turnpikes and Dissipativity

### v1

@@ 4.3 Turnpike leaving arc and linear end penalties
A well-known aspect of the turnpike property is the characteristic leaving arc, where state and input trajectories diverge from the turnpike in the final part of the considered time horizon, cf. Figure 1. While the formal Definition 4 does not require that a turnpike solution has a leaving arc, the early works [Dorfman et al., 1958, McKenzie, 1976] and many later results [Wilde and Kokotovic, 1972, Trélat and Zuazua, 2015] focus on it.
In the context of optimal control, the presence of a leaving arc arises when the values of the states $x$ and the adjoints $\lambda$ close to the turnpike differs substantially from their values at the end of the horizon. In optimal control there are three mechanism which generate leaving arcs [Faulwasser and Grüne, 2022]:
@@
  • Terminal stage costs $V_{\mathrm{f}}(x)$ which (without any active constraints at $k=N$) imply that the transversality condition $\lambda_{N}=\nabla V_{\mathrm{f}}(x)$.
  • Or the presence of state constraints which render it cheap to let a system drift towards the constraint boundary at the end of the horizon.
Interestingly, the same leaving arc phenomenon can be observed in GNEPs, see, e.g., the supply chain example [Hall et al., 2024a]. Moreover, the optimality interpretation of GNEPs—and in particular the structure of the KKT conditions, which is very similar to classic optimal control—suggest that the same mechanisms generate a leaving arc.
Inspired by the OCP analysis in [Zanon and Faulwasser, 2018, Faulwasser and Zanon, 2018], the next result shows how to suppress leaving arcs.
@@
Consider the GNEP (2) and let Assumptions 23 hold. Let $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v},\nu_{s}^{v})$ be a KKT point solving (25). Consider the following two statements:
@@
    and
    $\nabla V_{\mathrm{f}}^{v}(x_{s})+g_{x}^{\top}\mu_{N}^{v}=\lambda_{s}^{v}.$
@@
As per Theorem 7, part (i,) the steady-state primal-dual tuple $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v},\nu_{s}^{v})$ from (25) is a steady state of (24a24c). Moreover, using $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v},\nu_{s}^{v})$ to construct an $N$-step constant trajectory also the complementary slackness and the dual feasibility conditions (24e-24f) are satisfied as they correspond point-wise in time to their counterparts in (25).
It remains to analyze (24d). For case (i), notice that the Lagrangian $\mathcal{L}^{v}(x,u,\lambda^{v},\mu^{v},\nu^{v})$ in (23) has an additional term
$V_{\mathrm{f}}^{v}(x_{N})$
which, by standard KKT analysis leads to the modified boundary condition
$x_{0}=\mathbf{x}\quad\text{and}\quad\lambda_{N}=\nabla V_{\mathrm{f}}^{v}(x_{N})+g_{x}^{\top}\mu_{N}^{v}.$ (24d-i)
Recall that the assertion sets $\mathbf{x}=x_{s}$. From the condition that $\nabla V_{\mathrm{f}}^{v}(x_{s})+g_{x}^{\top}\mu_{N}^{v}=\lambda_{s}^{v}$ it then immediately follows that the $N$-step constant trajectory staying at $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v},\nu_{s}^{v})$ satisfies $x_{0}=x_{s}$ and (24d-i).
For case (ii), observe that the Lagrangian $\mathcal{L}^{v}(x,u,\lambda^{v},\mu^{v},\nu^{v})$ now includes the additional term
$\sigma^{\top}(x_{N}-x_{s}),$
where $\sigma\in\mathbb{R}^{n}$ is the additional multiplier for the point-wise terminal constraint. Once more, standard KKT analysis leads to the modified boundary condition
$x_{0}=\mathbf{x}\quad\text{and}\quad x_{N}=x_{s}.$ (24d-ii)
@@
The preceding analysis has shown that the constant trajectory staying at $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v})$ satisfies the KKT conditions (24). Theorem 5 then directly gives the assertion. This finishes the proof.
A first connection between the KKT system and the role of the terminal constraint on the dual variable was drawn in [Hall et al., 2024b, Prop. 3] and the proof therein. Specifically, it was demonstrated that with a terminal constraint $x_{N}=x_{s}$ and additional assumptions on the cost, the steady-state GNEP $(x_{s},u_{s})$ is the unique fixed point of the finite-horizon GNEP-KKT system and the unique equilibrium point of a closed-loop system resulting from a receding-horizon implementation of the GNE.
@@
Let the conditions of Proposition 9 hold and suppose that $\mu_{s}^{v}=0$ for all $v\in\mathcal{V}$. Then
@@
Indeed, quite similarly to the optimal control setting [Zanon and Faulwasser, 2018], also for GNEPs (2), the linear end penalty $V_{\mathrm{f}}^{v}(x)=x^{\top}\lambda_{s}^{v}$ admits a primal and a dual interpretation:
  • It directly specifies the boundary or terminal constraint on the dual variable $\lambda^{v}_{N}$. Interestingly, the point-wise boundary condition on $\lambda^{v}_{N}$ is an optimality conditions but it does not imply any reachability requirement of the actual primal state dynamics $x^{+}=f(x,u)$.
@@
    This trick is called rotating the stage cost in economic MPC [Diehl et al., 2011, Angeli et al., 2012]. The linear rotation with $\lambda_{s}$ implies that the steady constraint in (10) is only weakly active. This geometric modification of the stage cost with $\lambda_{s}$ has been called gradient correction in economic MPC, cf. [Zanon and Faulwasser, 2018].
@@ 4.4 Learning the linear end penalty
One may ask about the pros and cons of the different approaches that allow suppressing the leaving arc of turnpike solutions. Indeed, the point-wise terminal constraint $x_{N}=x_{s}$ requires solving the steady-state GNEP (10) as does the linear end penalty $V_{\mathrm{f}}^{v}(x)=x^{\top}\lambda_{s}^{v}$. In settings where the cost function changes occasionally—e.g. due to rare parametric updates (prices, etc.)—it is desirable to avoid the extra computational effort. Also, the primal terminal constraint $x_{N}=x_{s}$ jeopardize feasibility in case of short horizons.
In the context of suppressing turnpike leaving arcs in OCPs, [Faulwasser and Zanon, 2018, Sec. 4] proposed an adaptive strategy for the terminal penalty.
Algorithm 1 summarizes how this strategy may be transferred to GNEPs. In Step 2 the GNEP (2) is solved considering the linear end penalty $V_{\mathrm{f}}^{v}(x)=x^{\top}p^{v}_{i}$, where the initialization $p_{0}=0$ is used. Then, based on our turnpike insights, for all $v\in\mathcal{V}$, an approximation of the turnpike dual variable $\lambda^{v}_{s}$ is obtained by $p_{i}^{v}=\lambda^{v}(\frac{N}{2})$ in Step 5. The GNEP is resolved using the updated end penalty $V_{\mathrm{f}}^{v}(x)=x^{\top}p^{v}_{i}$. The algorithm ends if either the change in $\|p_{i}^{v}-p_{i+1}^{v}\|$ becomes small or an iteration limit is reached. Put differently, the algorithm builds upon the observation that whenever the states and inputs are close to their turnpike values also the dual variables stay close to their turnpike values. It is worth remarking that each agent can learn its specific linear end penalty. One may even learn this end penalty for some agent while others still exhibit leaving arcs.
While the formal convergence analysis is beyond the scope of the present paper, the next section will present numerical results demonstrating the efficacy of the proposed approach.
Algorithm 1 Learning $\lambda_{s}^{v}$ [Faulwasser and Zanon, 2018]
@@
3:while $i\leq i_{max}$ OR $\Delta_{i}\leq\varepsilon$ do
@@ 5 Simulation Study
The parameter values are $A=1.5$, $B^{1}=1$, $B^{2}=2$, $R^{1,1}=R^{1,2}=4$, $R^{2,2}=R^{2,1}=5$ and state weights $Q^{1}=1,Q^{2}=2$. The reference state is chosen as $x^{\text{ref}}=0.3$. Assumptions 2 and 3 are fulfilled by construction.
@@
Since $\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ and the steady-state GNE in (10) are not necessarily unique, the terminal penalty method requires verifying that $\lambda_{s}$ corresponds to the global turnpike. Local turnpike phenomenon in OCPs have been studied in [Krügel et al., 2023], but this is left for future work in the GNE setting.
@@ 6 Conclusion
In this manuscript we study GNE problems and derive system-theoretic insights into their open-loop trajectories based on dissipativity analysis. We extend the close relation between dissipativity and turnpike behaviors known in optimal control to dynamic generalized Nash games. Our simulation results show that the linear end penalties designed based on the optimality conditions successfully suppress the leaving arc of the turnpike even in the coupled, non-cooperative, game-theoretic setting.

### v3

@@ 4.3 Turnpike leaving arc and linear end penalties
A well-known aspect of the turnpike property is the characteristic leaving arc, where state and input trajectories diverge from the turnpike in the final steps of the horizon, cf. Figure 1. While the Definition 4 does not require a leaving arc, the early works [20, 45] and later results [2, 54] focus on it.
In the context of optimal control, the presence of a leaving arc arises when the values of the states $x$ and the adjoints $\lambda$ close to the turnpike differs substantially from their values at the end of the horizon. In optimal control there are three mechanisms which generate leaving arcs [28]:
@@
  • Terminal stage costs $V_{\mathrm{f}}(x)$ which (without any active constraints at $k=N$) imply the transversality condition $\lambda_{N}=\nabla V_{\mathrm{f}}(x)$.
  • Or the presence of state constraints and a cost which renders it cheap to let a system drift towards the constraint boundary at the end of the horizon.
Interestingly, the same leaving arc phenomenon can be observed in GNEPs, see, e.g., the supply chain example [38]. Moreover, the optimality interpretation of GNEPs—and in particular the structure of the KKT conditions, which is very similar to classic optimal control—suggest that the same mechanisms generate a leaving arc.
We have established in Proposition 1 that under a strict dissipativity assumption the steady-state GNE in (11). Thus, it is desirable to enforce that agents not only converge to $(x_{s},u_{s})$ but remain there.
Inspired by the OCP analysis in [29], the next result shows how to suppress leaving arcs in GNE trajectories.
@@
Consider the GNEP (2) and let Assumptions 45 hold. Let $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v},\nu_{s}^{v})$ be a KKT point solving (31). Let one of the following two statements hold:
@@
    and $\nabla V_{\mathrm{f}}^{v}(x_{s})+g_{x}^{\top}\mu_{N}^{v}=\lambda_{s}^{v}$.
@@
As per Theorem 8, part (i) the steady-state primal-dual tuple $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v},\nu_{s}^{v})$ from (31) is a steady state of (30a30c). Moreover, using $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v},\nu_{s}^{v})$ to construct an $N$-step constant trajectory also the complementary slackness and the dual feasibility conditions (30e-30f) are satisfied as they correspond point-wise in time to their counterparts in (31).
It remains to analyze (30d). For case (i), notice that the Lagrangian $\mathcal{L}^{v}(x,u,\lambda^{v},\mu^{v},\nu^{v})$ in (29) has an additional term $V_{\mathrm{f}}^{v}(x_{N})$ which using standard KKT analysis leads to the modified boundary condition
$x_{0}=\mathbf{x}\quad\text{and}\quad\lambda_{N}=\nabla V_{\mathrm{f}}^{v}(x_{N})+g_{x}^{\top}\mu_{N}^{v}.$ (30d-i)
Recall that the assertion sets $\mathbf{x}=x_{s}$. From the condition that $\nabla V_{\mathrm{f}}^{v}(x_{s})+g_{x}^{\top}\mu_{N}^{v}=\lambda_{s}^{v}$ it then immediately follows that the $N$-step constant trajectory staying at $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v},\nu_{s}^{v})$ satisfies $x_{0}=x_{s}$ and (30d-i).
For case (ii), observe that the Lagrangian $\mathcal{L}^{v}$now includes the additional term $\sigma^{\top}(x_{N}-x_{s})$, where $\sigma\in\mathbb{R}^{n}$ is the additional multiplier for the point-wise terminal constraint. Once more, standard KKT analysis leads to the modified boundary condition
$x_{0}=\mathbf{x}\quad\text{and}\quad x_{N}=x_{s}.$ (30d-ii)
@@
The preceding analysis has shown that the constant trajectory staying at $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v},\nu_{s}^{v})$ satisfies the KKT conditions (30). Theorem 6 then directly gives the assertion. This finishes the proof.
A first connection between the KKT system and the role of the terminal constraint on the dual variable was drawn in [35, Prop. 3] and the proof therein. Specifically, it was demonstrated that with a terminal constraint $x_{N}=x_{s}$ and additional assumptions on the cost, the steady-state GNEP $(x_{s},u_{s})$ is the unique fixed point of the finite-horizon GNEP-KKT system and the unique equilibrium point of a closed-loop system resulting from a receding-horizon implementation of the GNE.
Next, we design a linear end penalty which suppresses the leaving arc.
@@
Let the conditions of Proposition 10 (i) hold and suppose that $\mu_{s}^{v}=0$ for all $v\in\mathcal{V}$. Then
@@
Indeed, quite similarly to the optimal control setting [60], also for GNEPs (2), the linear end penalty $V_{\mathrm{f}}^{v}(x)=x^{\top}\lambda_{s}^{v}$ admits a primal and a dual interpretation:
  • It directly specifies the boundary or terminal constraint on the dual variable $\lambda^{v}_{N}$. Interestingly, the point-wise boundary condition on $\lambda^{v}_{N}$ is an optimality condition but it does not imply any reachability requirement of the actual primal state dynamics $x^{+}=f(x,u)$.
@@
    This is called rotating the stage cost in economic MPC [19, 3]. The linear rotation with $\lambda_{s}$ implies that the steady-state constraint in (11) is only weakly active. It is also called gradient correction, cf. [60].
@@ 4.4 Learning the linear end penalty
Both the point-wise terminal constraint $x_{N}=x_{s}$ and the linear end penalty approach $V_{\mathrm{f}}^{v}(x)=x^{\top}\lambda_{s}^{v}$ require solving the steady-state GNEP (11) in advance. Yet, in settings where the cost function changes occasionally (e.g. due to parametric updates such as prices) it would be desirable to avoid the extra computational effort. In the context of suppressing turnpike leaving arcs in OCPs, [29, Sec. 4] proposed an adaptive strategy for the terminal penalty.
Algorithm 1 summarizes how this strategy can be transferred to GNEPs. In Step 2 the GNEP (2) is solved considering the linear end penalty $V_{\mathrm{f}}^{v}(x)=x^{\top}p^{v}_{i}$, where the initialization $p_{0}=0$ is used. Then, based on our turnpike insights, for all $v\in\mathcal{V}$, an approximation of the turnpike dual variable $\lambda^{v}_{s}$ is obtained by $p_{i}^{v}=\lambda^{v}(\frac{N}{2})$ in Step 5. The GNEP is recomputed using the updated end penalty $V_{\mathrm{f}}^{v}(x)=x^{\top}p^{v}_{i}$. The algorithm ends if either the change in $\|p_{i}^{v}-p_{i-1}^{v}\|$ becomes small or an iteration limit is reached. Put differently, the algorithm builds upon the observation that whenever the states and inputs are close to their turnpike values also the dual variables stay close to their turnpike values. It is worth noting that each agent can learn its specific linear end penalty. One may even learn this end penalty for some agent while others still exhibit leaving arcs.
While the formal convergence analysis of Algorithm 1 is beyond the scope of this work , the next section will present numerical results demonstrating its efficacy.
Algorithm 1 Learning the linear end penalty $\lambda_{s}^{v}$
@@
3:while $i\leq i_{max}$ AND $\Delta_{i}\geq\varepsilon$ do
@@ 5 Simulation Study
The parameter values are $A=1.5$, $B^{1}=1$, $B^{2}=2$, $R^{1,1}=R^{1,2}=4$, $R^{2,2}=R^{2,1}=5$ and state weights $Q^{1}=1,Q^{2}=2$. The reference state is chosen as $x^{\text{ref}}=0.3$. Assumptions 4 and 5 are fulfilled by construction.
@@
Since $\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ and the steady-state GNE in (11) are not necessarily unique, the terminal penalty method requires verifying that $\lambda_{s}$ corresponds to the global turnpike. Local turnpike phenomenon in OCPs have been studied in [40], but this is left for future work in the GNE setting.
@@ 6 Conclusion
In this manuscript we study GNE problems and derive system-theoretic insights into their open-loop trajectories based on dissipativity analysis. We extend the close relation between dissipativity and turnpike behaviors known in optimal control to dynamic generalized Nash games. The two main ingredients of our analysis are a novel definition of a strict dissipativity property tailored to GNEPs as well as an assumption upper bounding the price of anarchy. Moreover, we establish a local variational characterization of the game value function and relate them to the geometry of storage functions.
Our simulation results show that the linear end penalties designed based on the optimality conditions successfully suppress the leaving arc of the turnpike even in the coupled, non-cooperative, game-theoretic setting.
@@
Appendix A Proof of Theorem 1
We adapt the original proof of [58] to the setting of dynamic GNEPs and to the strict dissipation inequality (14). Henceforth, we use the shorthand $\alpha_{\ell}(\cdot):=\alpha_{\ell}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)$. Where for simplicity we dropped the $^{*}$ notation.
[…]
$\Lambda_{\alpha_{\ell}}(\mathbf{x})\geq\sum_{k=0}^{m}\alpha_{\ell}\left(\cdot\right)-s(x_{k},u_{k})+\Lambda_{\alpha_{\ell}}(x_{m}).$ (38)
If at $\mathbf{x}=x_{0}$, we have $N^{*}<m$, optimality likewise implies the above inequality. Rearranging (38) directly gives (14).
[…]
$\Lambda(\mathbf{x})+\sum_{k=0}^{m}-\alpha_{\ell}\left(\cdot\right)+s(x_{k},u_{k})\geq\Lambda(x_{m})\geq-c>-\infty.$
The right hand side inequalities follow from the lower boundedness of storage functions $\Lambda$. Notice that these bounds hold for all $(x,u)\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$. Hence, we set $m=N$ and rearrange the inequalities such that
$\Lambda(\mathbf{x})\geq-c+\sup_{\begin{subarray}{c}N\in\mathbb{N}\\ (x,u)\\ \in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})\end{subarray}}-\sum_{k=0}^{N}-\alpha_{\ell}\left(\cdot\right)+s(x_{k},u_{k}).$
Here we have used that c-c is not affected by the optimization over the free end time $N$. With $x_{0}=\mathbf{x}$, the previous inequality is equivalent to $\Lambda(\mathbf{x})\geq-c+\Lambda_{\alpha_{\ell}}(\mathbf{x})$. Since $c\in\mathbb{R}$, $\Lambda_{\alpha_{\ell}}(\mathbf{x})=\infty$ would contradict $\Lambda(\mathbf{x})\in\mathbb{R}$. This finishes the proof.∎
Appendix B Performance bounds from bounded PoA
Lemma 12.
Suppose that $0<\nu\leq V^{\diamond}_{N}(\mathbf{x})\leq V<\infty$ is satisfied.
  • (i)
    If Assumption 2 holds with $P\in\mathbb{R}$, then
    $\sup_{(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})}J_{N}(x^{*},u^{*})-V^{\diamond}_{N}(\mathbf{x})\leq VP.$ (39)
  • (ii)
    If (39) holds with bound VPVP, then Assumption 2 holds with $\frac{V}{\nu}P+1$.
Proof B.1.
Part (i):  If Assumption 2 holds, we have
$\dfrac{1}{V^{\diamond}_{N}(\mathbf{x})}\displaystyle\sup_{(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})}J_{N}(x^{*},u^{*})\leq P\leq P+1.$
Multiplying both sides with $V^{\diamond}_{N}(\mathbf{x})\geq\nu>0$ gives
$\sup_{(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})}J_{N}(x^{*},u^{*})\leq V^{\diamond}_{N}(\mathbf{x})(P+1).$
Substracting $V^{\diamond}_{N}(\mathbf{x})$ and using $P\cdot V^{\diamond}_{N}(\mathbf{x})\leq PV$ gives the assertion. Part (ii):  Consider (39), add $V^{\diamond}_{N}(\mathbf{x})$ on both sides, and divide by $V^{\diamond}_{N}(\mathbf{x})$ to obtain
$\dfrac{1}{V^{\diamond}_{N}(\mathbf{x})}\displaystyle\sup_{(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})}J_{N}(x^{*},u^{*})\leq\dfrac{V}{V^{\diamond}_{N}(\mathbf{x})}P+1$
The left hand side is bounded from above by $\frac{V}{\nu}P+1$ which finishes the proof.